OpenStudy (anonymous):

Differentiate y according to x: x^2+x+xy =1

OpenStudy (anonymous):

you have to do implicit differentiation 2x + 1 + x.(dy/dx) + y = 0 then rearrange to get dy/dx =

OpenStudy (anonymous):

2x+1+xy'+y=0

OpenStudy (anonymous):

how can you make it into a fraction? As in: -(2x+y+1)/x

OpenStudy (anonymous):

if you NEED to get y' ONLY in terms of x, the maybe just rearrange the original in the 1st place: y = (1 - x - x^2)/x = (1/x) - 1 - x so that y' = -x^(-2) - 1

OpenStudy (anonymous):

That makes sense now! Thank you!

OpenStudy (anonymous):

you're welcome